OpenStudy (anonymous):

the equation of the line passing through the point (6, -2) and parallel to y = x + 7. is?

OpenStudy (anonymous):

For two lines to be parallel, they must have the same slope. In this case, the slope will be 1. The equation we will want to use is: \[y-y_1 = m(x-x_1)\] where m is the slope and x_1 and y_1 are the coordinates of any point on the line. We have a point and the slope, so we can plug in: \[y+2=1(x-6)\] Simplifying we find: \[y=x-8\]